### v1

# KARMA: Efficient Structural Defect Segmentation via Kolmogorov-Arnold Representation Learning

Abstract
Semantic segmentation of structural defects in civil infrastructure remains challenging due to variable defect appearances, harsh imaging conditions, and significant class imbalance. Current deep learning methods, despite their effectiveness, typically require millions of parameters, rendering them impractical for real-time inspection systems. We introduce KARMA (Kolmogorov-Arnold Representation Mapping Architecture), a highly efficient semantic segmentation framework that models complex defect patterns through compositions of one-dimensional functions rather than conventional convolutions. KARMA features three technical innovations: (1) a parameter-efficient Tiny Kolmogorov-Arnold Network (TiKAN) module leveraging low-rank factorization for KAN-based feature transformation; (2) an optimized feature pyramid structure with separable convolutions for multi-scale defect analysis; and (3) a static-dynamic prototype mechanism that enhances feature representation for imbalanced classes. Extensive experiments on benchmark infrastructure inspection datasets demonstrate that KARMA achieves competitive or superior mean IoU performance compared to state-of-the-art approaches, while using significantly fewer parameters (0.959M vs. 31.04M, a 97% reduction). Operating at 0.264 GFLOPS, KARMA maintains inference speeds suitable for real-time deployment, enabling practical automated infrastructure inspection systems without compromising accuracy. The source code can be accessed at the following URL: https://github.com/faeyelab/karma.
Index Terms:
Structural defect segmentation, Kolmogorov-Arnold networks, semantic segmentation, infrastructure inspection, parameter efficiency, real-time deployment.
I Introduction
Automated detection of structural defects in civil infrastructure is essential for public safety and efficient maintenance [1, 2]. The aging global infrastructure makes accurate segmentation of defects such as cracks, fractures, and deformations increasingly critical [3, 4]. Traditional methods relying on manual video inspection are labor intensive, slow, and error-prone [5]. In municipal sewer and culvert systems, missed defects can lead to catastrophic failures, compromising urban safety and water management. Although semantic segmentation methods offer pixel-level defect localization, existing approaches often struggle to balance accuracy with computational efficiency [3, 5].
The accurate segmentation of structural defects presents several challenges [6, 7]. Defects vary widely in size, shape, and appearance, occurring under various environmental conditions [8, 9]. Inspection imagery often suffers from poor lighting, debris obstructions, and inconsistent camera viewpoints [5]. Furthermore, there is a class imbalance, as rare defects such as holes or collapses occur far less frequently than common problems such as surface cracks [10]. Thus, robust and computationally efficient semantic segmentation methods are required for reliable performance in practical scenarios [6, 10].
Deep learning approaches, particularly Fully Convolutional Networks (FCNs) [11] and U-Net-based models [12, 13], have significantly improved segmentation performance through pixel-level predictions and innovations such as skip connections and feature pyramid modules. However, these models frequently involve millions of parameters, which limits deployment in resource-constrained settings such as inspection robots or real-time monitoring systems [14]. Therefore, there is a clear need for segmentation architectures that achieve high accuracy while remaining computationally efficient.
Kolmogorov–Arnold representation learning offers a promising solution to these challenges [15, 16, 17]. According to the Kolmogorov–Arnold theorem, any continuous multivariate function can be represented by compositions of one-dimensional functions, enabling compact and efficient neural network designs known as Kolmogorov-Arnold Networks (KANs) [18, 19, 20]. Recent applications of KAN in medical image segmentation achieved state-of-the-art results [21], highlighting their potential for structural defect segmentation.
We propose KARMA (Kolmogorov-Arnold Representation Mapping Architecture), a highly efficient semantic segmentation framework for structural defect detection. KARMA combines Kolmogorov-Arnold Networks (KANs) and Feature Pyramid Networks (FPNs) [22] in a novel way to tackle efficient segmentation challenges. Lightweight TiKAN modules within an adaptive FPN (AFPN) backbone use low-rank factorization for parameter efficiency and real-time speeds without losing accuracy. This tailored architecture meets specific problem needs, optimizing efficiency-performance trade-offs and outperforming current methods, as shown by extensive experimental results.
Our main contributions include:
  • A novel semantic segmentation architecture, KARMA, which uniquely integrates Kolmogorov-Arnold representation learning (specifically, parameter-efficient TiKAN modules) within an optimized encoder-decoder framework, representing the first problem-driven architectural synthesis of KAN-based methods for efficient structural defect segmentation.
  • Comprehensive empirical validation demonstrating competitive or superior performance compared to existing segmentation models, achieving state-of-the-art mean Intersection over Union (IoU) with significantly reduced parameters (up to 97% fewer) on challenging culvert and sewer defect datasets (CSDD) and S2DS.
  • An efficient inference pipeline suitable for practical deployment, accompanied by detailed analysis of speed-accuracy trade-offs and memory requirements for edge devices.
The remainder of this paper is structured as follows: Section II reviews related work in semantic segmentation architectures, structural defect detection, and Kolmogorov-Arnold networks. Section III details the theoretical background and architectural components of KARMA. Section IV outlines the experimental methodology and evaluation metrics. Section V presents results and comparative analysis. Section VI discusses broader implications and future research opportunities, concluding with final remarks in Section VII.
II Related Work
II-A Semantic Segmentation Architectures
Semantic segmentation has evolved significantly through innovations such as FCNs [11], enabling pixel-level predictions via end-to-end learning. U-Net [12] introduced an encoder-decoder architecture with skip connections, which enhances spatial and semantic information integration, notably in biomedical segmentation. Extensions such as UNet++ [23] and UNet 3+ [24] provided nested and dense skip connections, respectively, further improving multi-scale feature fusion. FPN [22] introduced top-down pathways, widely adopted for segmentation tasks, while BiFPN [25] incorporated bidirectional connections and learned feature weighting for improved efficiency. Recent architectures, including EGE-UNet [26] and SA-UNet [27], use edge guidance and attention mechanisms to improve boundary accuracy with fewer parameters. Transformer-based approaches like Swin-UNet [28] and Segformer [29] use attention mechanisms to capture global dependencies efficiently. Lightweight models such as Rolling UNet [30] and MobileUNETR [31] exemplify high accuracy combined with reduced computational demands, aligning with the efficiency objectives of KARMA.
II-B Structural Defect Segmentation
Recent developments in structural defect segmentation have shifted from classification and object detection techniques towards precise pixel-level segmentation methods [6, 32]. Improved architectures based on U-Net have significantly enhanced accuracy in clearly defining defect shapes [5]. To overcome issues such as class imbalance, methods like Sewer-ML and enhanced FPN (E-FPN) emphasize multi-scale defect representation and customized loss functions [33, 34]. Leading-edge approaches integrate Haar-like features and optimized pyramid structures to achieve high segmentation accuracy [35]. Recent optimization strategies, including focal loss, class-weighted methods, and boundary-aware loss functions, further improve segmentation performance, particularly for uncommon defect classes. Current research trends prioritize achieving both accuracy and computational efficiency by employing pyramid structures and attention mechanisms suitable for real-time applications [6, 32].
II-C Kolmogorov-Arnold Networks and Representation Learning
Kolmogorov-Arnold Network (KAN) [18], inspired by the Kolmogorov-Arnold representation theorem, provides compact, but powerful representations by decomposing complex functions into compositions of one-dimensional functions. Recent applications of KANs in medical imaging and defect detection highlight their potential to achieve high accuracy [21]. Hybrid architectures such as HKAN [36] integrate CNNs and transformers with KAN layers to effectively handle complex data patterns. KARMA uniquely incorporates tiny KAN principles within an FPN-style architecture, specifically targeting efficient and accurate segmentation of structural defects, representing the first such integration in this domain.
II-D Bridging the Gap: From Existing Methods to KARMA
Despite progress in the literature, a gap remains in achieving both top-tier segmentation accuracy and extreme computational efficiency, essential for real-time, on-device defect inspection. High-accuracy models like advanced U-Net variants and Transformer-based architectures incur high computational costs, limiting applicability. Lightweight models, lacking performance, struggle with complex defect patterns. Though KANs are parameter-efficient compared to MLPs, their use in complex frameworks like U-KAN is not fully optimized, leading to heavy models. Our work, KARMA, addresses this efficiency-performance gap by using optimized, low-rank TiKAN modules in a custom feature pyramid network, reimagining efficient and accurate defect segmentation. It overcomes previous limitations, showing that Kolmogorov-Arnold principles can create models that are vastly more efficient while matching or exceeding current accuracy standards.
III Problem Formulation
Let $\mathcal{I}=\{I_{i}\}_{i=1}^{N}$ be a set of $N$ input images, where each image $I_{i}\in\mathbb{R}^{H\times W\times C}$ has height $H$, width $W$, and $C$ channels. For each image $I_{i}$, there exists a corresponding ground-truth segmentation mask $M_{i}\in\{0,1\}^{H\times W\times K}$, where $K$ denotes the number of defect categories including the background class. The objective is to learn a mapping function $f_{\theta}:\mathcal{I}\rightarrow\mathcal{M}$ parameterized by $\theta$ that can effectively segment structural defects.
Given an adaptive feature pyramid network (AFPN) backbone $\phi_{AFPN}$ that extracts multi-scale features $\{F_{l}\}_{l=1}^{L}$ at $L$ different scales, where $F_{l}\in\mathbb{R}^{H_{l}\times W_{l}\times C_{l}}$, our goal is to incorporate Kolmogorov-Arnold (KA) representation learning to improve the feature learning process. The term ‘adaptive’ in our adaptive feature pyramid network (AFPN) refers to its enhanced functional capability to dynamically adjust feature representations. This adaptivity is primarily achieved through the strategic integration of our TiKAN modules (detailed in Section IV-A). As will be elaborated, these TiKAN modules possess learnable scaling factors (e.g., $s_{\text{base}}$ and $s_{\text{spline}}$ in Equation 6) that allow the network to dynamically balance linear base and non-linear spline transformations for features at each FPN level. Furthermore, the KANBlock components within our TiKAN Enhancement Module (Section IV-B-2) utilize residual learning, enabling them to adaptively determine the degree of feature modification. Thus, our AFPN is not presented as a structurally novel FPN in terms of its fundamental pyramidal connections, but rather as an FPN framework rendered functionally adaptive by these integrated KAN-based components, allowing it to better respond to the diverse characteristics of input features across different scales. For each feature level $l$, we define a TiKAN module $\psi_{l}$ that transforms the features as:
$\hat{F}_{l}=\psi_{l}(F_{l};\Theta_{l}),$ (1)
where $\Theta_{l}$ represents the learnable parameters of the TiKAN module at level $l$. Following the Kolmogorov-Arnold representation theorem, $\psi_{l}$ is designed to approximate any continuous multivariate function through a composition of continuous univariate functions and basic arithmetic operations.
The complete KARMA architecture can be formulated as:
$\hat{M}=f_{\theta}(I)=\sigma(\mathcal{D}(\{\psi_{l}(F_{l};\Theta_{l})\}_{l=1}^{L})),$ (2)
where $\mathcal{D}$ represents the decoder module that fuses the TiKAN-enhanced multi-scale features to produce the final segmentation, and $\sigma$ is the softmax activation function that normalizes the network outputs into class probabilities.
IV Proposed Kolmogorov-Arnold Representational Mapping Architecture (KARMA)
We propose KARMA, a novel defect segmentation architecture that enhances the AFPN backbone with Kolmogorov-Arnold representation learning through lightweight TiKAN modules as shown in Fig. 1. Unlike U-KAN which utilizes the original KAN layers in UNet, our method incorporates an efficient TiKAN architecture into an AFPN framework. The key innovation lies in our low-rank formulation of the KA representation that significantly reduces parameters while maintaining expressiveness. KARMA consists of three main components: (1) An AFPN backbone for hierarchical feature extraction; (2) Low-rank TiKAN modules that transform features at each scale using efficient KA representation learning; and (3) A decoder that fuses the enhanced multi-scale features for final segmentation.
Figure 1: KARMA architecture overview showing the three main components: bottom-up pathway with InceptionSepConv blocks (c1-c5), TiKAN enhancement module at the deepest level (c5), and top-down pathway with feature fusion (p2-p5)
IV-A TiKAN: Low-Rank KA Representation Learning
Our TiKAN design is motivated by the Kolmogorov-Arnold (K-A) representation theorem, which states that any continuous multivariate function $f:\mathbb{R}^{n}\to\mathbb{R}$ can be decomposed into univariate functions ${\psi_{p}}_{p=1}^{n}$ and ${\Phi_{q}}_{q=0}^{2n}$:
$f(x_{1},x_{2},\dots,x_{n})=\sum_{q=0}^{2n}\Phi_{q}\left(\sum_{p=1}^{n}\psi_{p}(x_{p})+c_{q}\right),$ (3)
where $c_{q}$ are constants. This fundamental theorem suggests that complex high-dimensional functions can be represented through compositions of simpler one-dimensional functions. Building on this theoretical foundation, we design TiKAN to learn complex feature transformations through compositions of one-dimensional functions while maintaining computational efficiency through low-rank decomposition.
IV-A1 Low-Rank Base Transformation
Given an input feature map $F_{l}\in\mathbb{R}^{H_{l}\times W_{l}\times C_{l}}$ at level $l$, TiKAN first applies a low-rank base transformation:
$\phi_{base}(F_{l})=\sigma(W_{u}^{T}W_{v}F_{l}+b),$ (4)
where $W_{u}\in\mathbb{R}^{C_{l}\times r}$ and $W_{v}\in\mathbb{R}^{r\times C_{l}}$ represent the low-rank decomposition with rank $r$, $b$ is the bias term, and $\sigma$ is the SiLU activation function. This base transformation serves as the primary feature mapping mechanism, while the low-rank decomposition significantly reduces the parameter count compared to a standard linear transformation. The SiLU activation introduces non-linearity while maintaining smooth gradients, which is crucial for learning complex feature relationships.
IV-A2 Spline-Based Nonlinear Transformation
Following the KA theorem’s principle of univariate function composition, we implement the spline transformation using learnable one-dimensional functions:
$\phi_{spline}(F_{l})=S_{u}^{T}S_{v}g(F_{l}),$ (5)
where $S_{u}\in\mathbb{R}^{C_{l}\times r_{f}}$ and $S_{v}\in\mathbb{R}^{r_{f}\times(G+O)}$ form a factorized spline weight matrix. The parameter $r_{f}$ represents the factor rank for spline weights, which controls the expressiveness of the spline transformation. The grid size $G$ determines the granularity of the spline approximation, while the spline order $O$ defines the smoothness of the interpolation. The function $g(.)$ generates grid points with optional noise regularization during the training process.
The spline transformation provides additional flexibility in modeling nonlinear relationships. The grid-based approach allows for precise control over the function approximation, while the noise regularization during training helps prevent overfitting and improves generalization. The factorized spline weight matrix maintains parameter efficiency while allowing for rich representational capacity.
The complete TiKAN transformation combines these components with learnable scaling factors:
$\psi_{l}(F_{l})=\phi_{base}(F_{l})\cdot s_{base}+\phi_{spline}(F_{l})\cdot s_{spline}.$ (6)
The scaling factors $s_{base}$ and $s_{spline}$ allow the network to adaptively balance the contributions of the base and spline transformations. This adaptive mechanism is particularly important as different feature levels may require different degrees of nonlinear transformation. This equation (Equation 6) represents the overarching principle of our TiKAN approach, which aims to achieve parameter-efficient feature transformation by combining a low-rank base transformation and a learnable spline-based non-linear transformation. The specific KAN-inspired components that realize this principle, namely KANLinear, KANLayer, and KANBlock, are detailed in the subsequent sections, particularly in the context of our TiKAN Enhancement Module (Section IV-B-2).
IV-A3 Parameter Efficiency through Constraints
To maintain efficiency while preserving the expressive power of KA representation, we employ several key constraints. First, we utilize low-rank weight matrices that reduce parameters from $O(C_{l}^{2})$ to $O(C_{l}(r+r_{f}))$, making the model more computationally tractable without significantly sacrificing performance. Second, we implement shared spline functions across channels as a deliberate architectural design choice for parameter efficiency within our TiKAN modules. This means a common set of learnable spline parameters (or their defining factorized weights $S_{u},S_{v}$) is reused across multiple feature channels, rather than each channel-to-channel path having unique splines. The phrase “when appropriate” refers to this structural design for efficiency, utilizing the assumption that certain fundamental non-linear transformations can be effectively reused across different feature channels, significantly reducing parameters while maintaining expressive power. Third, we incorporate pruning of redundant connections during training through the following mechanism:
$W_{pruned}=W\odot\mathbbm{1}(|W|>\tau),$ (7)
where $\tau$ is a small threshold and $\mathbbm{1}(.)$ is the indicator function. This pruning mechanism helps eliminate weak connections and improve model sparsity.
In practice, we initialize the grid points uniformly in the range [1,1][-1,1] and set the spline order to 3 for cubic interpolation. The factor ranks $r$ and $r_{f}$ are chosen based on the channel dimension $C_{l}$ to maintain a good balance between model capacity and efficiency. During training, we employ gradient clipping to ensure stable optimization of the spline weights.
IV-B Integration of TiKAN into AFPN
Our KARMA architecture strategically integrates TiKAN modules into the AFPN backbone through a hierarchical feature pyramid design. The architecture consists of bottom-up and top-down pathways with TiKAN enhancement at critical stages. The first component addresses the challenge of efficient feature extraction through InceptionSepConv blocks, while the second component leverages TiKAN for enhanced representation learning. The multi-scale design allows the network to capture both fine-grained defect details and global contextual information.
IV-B1 Bottom-up Pathway with InceptionSepConv
Our bottom-up pathway architecture consists of five sequential stages, each designed to extract increasingly abstract feature representations. For an input image $x$, the pathway produces a hierarchical set of feature maps as follows:
$c_{i}=\begin{cases}\mathcal{F}_{1}(x),&\text{if }i=1,\\ \mathcal{F}_{i}\bigl{(}\mathcal{P}(c_{i-1})\bigr{)},&\text{if }i\in\{2,3,4,5\},\end{cases}$
where $\mathcal{F}_{i}=\text{InceptionSepConv}_{i}$ and $\mathcal{P}=\text{Pool}$ represent the InceptionSepConv blocks and pooling operations, respectively. Each InceptionSepConv block incorporates three parallel branches optimized for multi-scale feature extraction. The first branch utilizes two depthwise-separable $3\times 3$ convolutions that efficiently extract local features. The second branch employs two depthwise-separable $5\times 5$ convolutions that capture broader contextual information. The third branch consists of a MaxPool operation followed by a $1\times 1$ convolution to compress spatial information while preserving essential features. To facilitate comprehensive feature representation while maintaining computational efficiency, we implement a progressive expansion of channel dimensions across the network ($3\rightarrow 48\rightarrow 96\rightarrow 192\rightarrow 384\rightarrow 576$). This systematic increase in dimensionality enables the extraction of increasingly complex features through the network hierarchy while effectively managing computational resources. The architecture thus achieves an optimal balance between representational capacity and operational efficiency.
IV-B2 TiKAN Enhancement Module
At the deepest level of feature hierarchy ($c_{5}\in\mathbb{R}^{B\times 576\times 8\times 8}$), which is the output of the final stage of the bottom-up pathway as depicted in Fig. 1, we implement a TiKAN enhancement module that transforms the feature representation through Kolmogorov-Arnold learning principles:
$\mathbf{O}=(\mathcal{R}\circ\mathcal{K}\circ\mathcal{E})(c_{5}),$ (8)
where $\mathcal{E}$, $\mathcal{K}$, and $\mathcal{R}$ represent the PatchEmbed, KANBlock, and Unpatchify operations, respectively. To clarify the hierarchy of KAN-based components used within our TiKAN Enhancement Module, we define them as follows: The most fundamental unit is the KANLinear operation, which provides the core KAN-like learnable activation functionality. Building upon this, a KANLayer in our architecture is a more complex structure composed of KANLinear operations and other layers. Finally, a KANBlock encapsulates a KANLayer with additional standard neural network operations like normalization and residual connections. We detail these components in the context of the TiKAN Enhancement Module below:
The PatchEmbed operation ($\mathcal{E}$) transforms the feature map into a sequence of tokens, enabling more flexible feature manipulation. Each token captures a local spatial region in the feature space, allowing the network to model long-range dependencies more efficiently.
The KANBlock operation ($\mathcal{K}$) is a higher-level module that wraps our KANLayer (defined next) with standard architectural elements. It consists of: (1) input normalization via LayerNorm; (2) the KANLayer transformation (which internally uses KANLinear operations and depthwise separable convolutions with activations); and (3) a residual connection to facilitate gradient flow and adaptive feature modification.
The KANLayer transformation is a composite structure that utilizes the fundamental KANLinear operations. Specifically, our KANLayer consists of two sequential KANLinear instances, with each KANLinear operation followed by a depthwise convolution, BatchNorm, and ReLU (denoted as a ‘DW_bn_relu‘ module in our implementation) to facilitate spatial mixing of token features.
The KANLinear operation (denoted as $\mathcal{K}_{L}$ and conceptually represented by Equation 9) is the foundational KAN-inspired building block in our model. It replaces a standard linear transformation and fixed activation with a layer where learnable spline-based activation functions ($\boldsymbol{\phi}_{s}$) are combined with a learnable base function ($\boldsymbol{\phi}_{b}$), each scaled by learnable factors ($\mathbf{s}_{b},\mathbf{s}_{s}$). Internally, it employs a low-rank linear transformation for its base component and factorized weights for its spline components to ensure parameter efficiency.
$\mathcal{K}_{L}(\mathbf{x})=\boldsymbol{\phi}_{b}(\mathbf{x})\cdot\mathbf{s}_{b}+\boldsymbol{\phi}_{s}(\mathbf{x})\cdot\mathbf{s}_{s},$ (9)
where $\mathcal{K}_{L}$ denotes the KANLinear operation, $\boldsymbol{\phi}_{b}$ and $\boldsymbol{\phi}_{s}$ represent the base and spline activation functions, and $\mathbf{s}_{b}$ and $\mathbf{s}_{s}$ are their respective scaling factors.
The enhanced feature representation $\mathbf{O}$ is subsequently projected to the required dimensionality through a dimensionality reduction operation, defined as $\mathbf{p}_{5}=\mathcal{C}(\mathbf{O})$, where $\mathcal{C}$ represents a $1\times 1$ convolutional operation that reduces the channel dimensionality while preserving spatial information. Our experiments demonstrate that this TiKAN enhancement module significantly improves the network’s ability to capture complex defect patterns while adding only minimal computational overhead, achieving a favorable balance between computational efficiency and representation power.
IV-B3 Top-down Pathway with Feature Fusion
The top-down pathway implements a progressive feature fusion strategy through depthwise-separable convolutions and upsampling operations, expressed as $\mathbf{p}_{i}=\mathcal{D}(\mathbf{c}_{i})+\mathcal{U}_{2}(\mathbf{p}_{i+1})$ for $i\in\{4,3,2\}$, where $\mathcal{D}$ represents a depthwise-separable convolution with kernel size $1$, and $\mathcal{U}_{2}$ denotes a $2\times$ upsampling operation. This design choice significantly reduces computational cost while maintaining effective feature transformation capability. The additive connections allow the network to combine high-level semantic information from deeper layers with lower-level spatial details from shallower layers, creating a rich multi-scale representation that preserves both fine-grained features and broader contextual information.
IV-B4 Multi-scale Prediction Heads
At each scale level in the feature pyramid, we deploy specialized prediction heads implemented through carefully designed convolutional layers, expressed as $\mathbf{o}_{l}=\mathcal{C}_{3}(\mathbf{p}_{l})$ for $l\in\{2,3,4,5\}$, where $\mathcal{C}_{3}$ represents a $3\times 3$ convolutional operation applied to the feature map $\mathbf{p}_{l}$. These scale-specific predictions capture information at different granularities—from fine-grained details at lower levels to broader contextual patterns at higher levels. The multi-scale predictions are subsequently unified through an adaptive fusion mechanism: $\mathbf{o}_{f}=\sum_{l=2}^{5}\mathcal{U}_{2^{l}}(\mathbf{o}_{l})$, where $\mathcal{U}_{2^{l}}$ denotes upsampling by a factor of $2^{l}$ and $\mathbf{o}_{f}$ represents the final fused output. This fusion process employs progressive upsampling operations to align predictions from different scales, followed by element-wise addition to combine complementary information across the feature hierarchy.
IV-B5 Static Dynamic Prototype Mechanism
To address class imbalance and the need for precise defect distinction, KARMA includes a static-dynamic prototype mechanism. This mechanism improves the network’s ability to learn feature representations, especially for minority classes and ambiguous defect patterns.
The static part of this mechanism involves creating representative ‘prototypes’ for each defect class. These prototypes are learned feature vectors that represent the key characteristics and variations within each category. During training, the network adjusts its learned feature representations to align with these class-specific prototypes. This process helps stabilize the feature space, providing reference points that help the model form distinct clusters for each class, even for those with few training examples.
The dynamic part describes how these prototypes are used during network operation, in both training and inference. Instead of only feed-forward classification, the dynamic mechanism allows an adaptive interaction between incoming feature maps and the learned prototypes. This interaction can involve the network adjusting its feature processing based on similarity to the prototypes. For example, features from unclear regions can be made more distinct by relating them to their most relevant prototype, which increases their discriminative power. This dynamic adjustment is important for improving segmentation precision, especially at boundaries, and for consistent performance across all defect types, including rare or subtle ones.
Overall, this static-dynamic prototype mechanism supports a more effective learning process. By providing explicit or implicit reference points in the feature space and using them dynamically, KARMA achieves high segmentation accuracy and stability, especially in challenging real-world situations with significant class imbalance and diverse defect appearances.
IV-C Loss Functions and Training Strategy
IV-C1 Multi-component Loss Function
We propose a comprehensive loss function that combines three complementary components, expressed as $\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{ce}}+\beta\mathcal{L}_{\text{dice}}+\gamma\mathcal{L}_{\text{reg}}$, where the weighting coefficients $\alpha$, $\beta$, and $\gamma$ are empirically determined as 0.5, 0.3 and 0.2, respectively, balancing the contributions of each term.
The primary component, cross-entropy loss $\mathcal{L}_{\text{ce}}$, addresses the fundamental pixel-wise classification task:
$\mathcal{L}_{\text{ce}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}w_{k}M_{i,k}\log(\hat{M}_{i,k}).$ (10)
We introduce class-specific weights $w_{k}$ that dynamically adjust the learning focus based on class frequencies:
$w_{k}=\frac{\text{median}(f_{k})}{f_{k}},\quad f_{k}=\frac{\text{\# pixels of class k}}{\text{total \# pixels}}.$ (11)
This adaptive weighting scheme effectively handles the inherent class imbalance common in defect segmentation tasks.
To specifically target boundary accuracy, we incorporate the Dice loss $\mathcal{L}_{\text{dice}}$:
$\mathcal{L}_{\text{dice}}=1-\frac{2\sum_{i=1}^{N}\sum_{k=1}^{K}M_{i,k}\hat{M}_{i,k}+\epsilon}{\sum_{i=1}^{N}\sum_{k=1}^{K}(M_{i,k}+\hat{M}_{i,k})+\epsilon}.$ (12)
The Dice loss provides a more geometrically meaningful measure of segmentation quality, particularly effective at improving boundary delineation.
IV-C2 Regularization Components
Our regularization framework incorporates multiple complementary terms combined into a unified regularization loss $\mathcal{L}_{\text{reg}}$:
$\mathcal{L}_{\text{reg}}=\lambda_{1}\mathcal{L}_{\text{smooth}}+\lambda_{2}\mathcal{L}_{\text{sparsity}},$ (13)
where $\lambda_{1}$ and $\lambda_{2}$ are weighting coefficients empirically set to 0.1 and 0.01 respectively.
The smoothness component $\mathcal{L}_{\text{smooth}}$ ensures the learned KAN functions exhibit desirable continuity properties:
$\mathcal{L}_{\text{smooth}}=\sum_{l=1}^{L}|\nabla^{2}\psi_{l}|^{2}+|\nabla^{2}\phi_{l}|^{2}.$ (14)
This term computes the second-order derivatives of both the inner ($\psi_{l}$) and outer ($\phi_{l}$) learned functions across all L layers of the network. By minimizing these second-order variations, we encourage the network to learn smooth functional mappings that are less likely to overfit.
To complement the smoothness constraints, we introduce a sparsity-inducing regularization term:
$\mathcal{L}_{\text{sparsity}}=\sum_{l=1}^{L}|W_{l}|_{1}.$ (15)
This L1 regularization term on the network weights $W_{l}$ encourages parameter efficiency by driving unnecessary connections towards zero, effectively learning a more compact model.
IV-C3 Optimization Strategy
Our optimization strategy involves a single-stage training for 50 epochs using the AdamW optimizer (weight decay of 1e-5, initial learning rate 0.001) combined with a cosine annealing scheduler reducing the learning rate to 1e-6. Data is processed in batches of 16 with gradient clipping (maximum norm 1.0) for training stability. The loss function integrates weighted cross-entropy loss, Dice loss, and Focal loss (weights 0.5, 0.3, and 0.2 respectively), effectively managing class imbalance and ensuring precise boundary representation. The best-performing model is selected based on validation IoU scores to prevent overfitting. Training is conducted in PyTorch on an NVIDIA A100 GPU, ensuring stable convergence and high segmentation accuracy.
V Experimental Results and Analysis
To ensure a comprehensive and up-to-date evaluation, we compare KARMA against a wide array of state-of-the-art defect segmentation methods, including those recently proposed in 2023 and 2024. Our benchmarks encompass traditional CNN-based architectures, advanced U-Net variants, and prominent Transformer-based models, reflecting the latest advancements in the field.
V-A Experimental Setting
V-A1 Datasets
We assess our proposed method using two challenging structural defect datasets. The Structural Defects Dataset (S2DS) [37] contains 743 high-resolution images of concrete surfaces collected via DSLR cameras, mobile phones, and drones. These images are pixel-wise annotated across seven distinct classes: background, cracks (linear fractures), spalling (surface detachment), corrosion (rust), efflorescence (chemical deposits), vegetation (plant growth), and control points (fiducial markers). The dataset is divided into training (563 images), validation (87 images), and testing (93 images) subsets. S2DS is characterized by significant class imbalance (dominated by background pixels), diverse imaging conditions, and subtle textural variations among defects, providing a rigorous benchmark for evaluating model robustness and generalization. Additionally, the Culvert-Sewer Defects Dataset comprises around 6,300 frames [34] extracted from 580 annotated underground inspection videos supplied by the U.S. Army Corps of Engineers and an industry partner. This dataset is split into 70% training, 15% validation, and 15% testing subsets. Frames are manually annotated according to NASSCO PACP standards, covering eight defect classes: cracks, roots, holes, joint issues, deformation, fracture, encrustation/deposits, and loose gasket. These annotations represent real-world scenarios with substantial class imbalance, variable materials, challenging imaging conditions, and multi-scale defect features, making it ideal for testing KARMA’s effectiveness in realistic infrastructure inspection contexts.
V-A2 Benchmark Algorithms
KARMA is compared with 16 benchmark algorithms representing state-of-the-art approaches in image segmentation. These include traditional CNN-based FCN architectures: U-Net [12], FPN [22], Attention U-Net [38], UNet++ [23], BiFPN [25], SA-UNet [27], UNet3+ [24], UNeXt [39], EGE-UNet [26], and Rolling UNet [30]. The comparison also includes transformer-based segmentation models: HierarchicalViT U-Net [40], Swin-UNet [41], MobileUNETR [31], Segformer [29], and FasterVit [42]. Lastly, we compare with a KAN-based U-Net called U-KAN [21].
V-A3 Metrics
Our experimental evaluation was conducted on the CSDD dataset using standard segmentation performance metrics. We report parameters (in millions) and computational complexity (in GFLOPS) to measure model efficiency. For segmentation quality, we report F1 Score (with and without background), mean Intersection over Union (mIoU) (with and without background), Balanced Accuracy, Mean Matthews Correlation Coefficient, and Frequency Weighted IoU.
V-B KARMA Model Variants
To provide a comprehensive evaluation across different computational budgets and performance targets, we introduce three variants of our KARMA architecture: the baseline KARMA, KARMA High, and KARMA Flash. These variants differ primarily in their channel dimensions throughout the network, the complexity of the KAN module application, and the channel widths in the Feature Pyramid Network (FPN) stages. The motivation behind these variants is to explore the trade-offs between model capacity, computational efficiency (parameters, FLOPs, inference speed), and segmentation accuracy.
KARMA (Baseline): This is the standard configuration detailed in Section IV, designed to balance high accuracy and efficiency. Its key characteristics include a bottom-up pathway culminating in 576 channels for the $c_{5}$ feature map, with the KANBlock in the TiKAN Enhancement Module operating on these 576 channels. Additionally, the FPN layers in the top-down pathway operate with 64 channels, and within the KANLayer, the hidden feature dimension is equal to the input feature dimension.
KARMA High: This variant is designed to maximize segmentation performance, potentially at the cost of increased computational resources, with the aim of establishing an upper bound on the performance achievable with the KARMA architectural principles by scaling up its capacity. The main differences from the baseline KARMA include increased channel dimensions in the bottom-up InceptionSepConv blocks, resulting in the $c_{5}$ feature map having 1024 channels. Additionally, the KANBlock in the TiKAN Enhancement Module processes these 1024 channels, and the FPN layers in the top-down pathway are wider, operating with 128 channels and utilizing standard convolutions instead of depthwise-separable convolutions for potentially richer feature fusion.
KARMA Flash: This variant is designed for maximum computational efficiency and inference speed, ideal for environments with limited resources or applications requiring real-time processing. It aims to create a lightweight model while preserving the key benefits of the KARMA architecture through several modifications compared to the baseline. These modifications include reducing channel dimensions in the later stages of the bottom-up pathway, with the $c_{5}$ feature map reduced to 384 channels. A pre-KAN $1\times 1$ convolution further decreases the $c_{5}$ features from 384 to 256 channels before they are processed by the KANBlock. The KANBlock in the TiKAN Enhancement Module operates on these reduced 256 channels. Within the KANLayer, the hidden feature dimension is halved relative to its input feature dimension, further reducing parameters in the KAN operations, and the FPN layers in the top-down pathway are narrower, operating with only 32 channels.
These three variants allow us to demonstrate the scalability and flexibility of the KARMA design, showcasing its adaptability to different performance requirements and computational constraints.
TABLE I: Performance Metrics Comparison Across Different Models on CSDD
Model (Year) Params F1 Score mIoU Bal. Acc. Mean MCC FW IoU
(M) GFLOPS w/bg w/o w/bg w/o
U-Net[12] 31.04 13.69 0.853 0.836 0.756 0.729 0.855 0.838 0.761
FPN[22] 21.20 7.809 0.848 0.830 0.748 0.719 0.825 0.833 0.768
Att. U-Net[38] 31.40 13.97 0.860 0.843 0.765 0.738 0.865 0.845 0.773
UNet++[23] 4.984 6.462 0.847 0.829 0.747 0.719 0.808 0.832 0.761
BiFPN[25] 4.459 17.76 0.847 0.829 0.747 0.719 0.824 0.833 0.773
SA-UNet[27] 7.857 3.625 0.855 0.839 0.760 0.733 0.851 0.842 0.780
UNet3+[24] 25.59 33.04 0.849 0.832 0.751 0.722 0.866 0.835 0.780
UNeXt[39] 6.294 1.163 0.850 0.832 0.752 0.724 0.826 0.835 0.760
EGE-UNet[26] 3.025 0.306 0.808 0.786 0.695 0.660 0.771 0.791 0.706
Rolling UNet-L[30] 28.33 8.222 0.854 0.838 0.758 0.730 0.865 0.840 0.776
HierarchicalViT U-Net[40] 14.58 1.312 0.829 0.810 0.724 0.693 0.841 0.814 0.746
Swin-UNet[41] 14.50 0.983 0.835 0.816 0.732 0.702 0.837 0.816 0.728
MobileUNETR[31] 12.71 1.068 0.829 0.809 0.722 0.691 0.813 0.811 0.729
Segformer[29] 13.67 0.780 0.835 0.816 0.732 0.702 0.851 0.816 0.729
FasterVit[42] 25.23 1.571 0.823 0.802 0.717 0.686 0.836 0.804 0.723
U-KAN[21] 25.36 6.905 0.853 0.836 0.757 0.729 0.838 0.839 0.776
U-TiKAN 9.50 6.905 0.853 0.836 0.756 0.728 0.858 0.837 0.764
KARMA (this paper) 0.959 0.264 0.855 0.838 0.759 0.731 0.835 0.840 0.774
Note: bg = background, Bal. Acc. = Balanced Accuracy, FW = Frequency Weighted,
MCC = Matthews Correlation Coefficient
TABLE II: Performance Metrics Comparison Across Different Models on S2DS
Model (Year) Params F1 Score mIoU Bal. Acc. Mean MCC FW IoU
(M) GFLOPS w/bg w/o w/bg w/o
U-Net[12] 26.08 40.03 0.6792 0.6310 0.5744 0.5138 0.7483 0.6500 0.7862
FPN[22] 21.18 31.14 0.7126 0.6714 0.5851 0.5286 0.8152 0.6824 0.7418
Att. U-Net[38] 31.40 55.87 0.7376 0.7000 0.6127 0.5598 0.8384 0.7096 0.7648
UNet++[23] 4.983 25.83 0.7250 0.6850 0.6033 0.5485 0.7747 0.6976 0.7591
BiFPN[25] 4.458 68.41 0.8072 0.7784 0.6968 0.6527 0.8425 0.7837 0.8554
SA-UNet[27] 7.855 14.49 0.6951 0.6490 0.5987 0.5411 0.7519 0.6660 0.8059
UNet3+[24] 25.59 132.0 0.7240 0.6837 0.6047 0.5498 0.7789 0.6978 0.7678
UNeXt[39] 6.293 4.641 0.7557 0.7184 0.6372 0.5835 0.8207 0.7344 0.8529
EGE-UNet[26] 2.832 3.550 0.6320 0.5757 0.5230 0.4533 0.6162 0.5891 0.7789
Rolling UNet-L[30] 28.33 32.88 0.7347 0.6949 0.6181 0.5630 0.7789 0.7061 0.8129
HierarchicalViT U-Net[40] 14.77 5.241 0.6705 0.6207 0.5636 0.5008 0.7278 0.6448 0.7808
Swin-UNet[41] 2.632 1.321 0.6166 0.5570 0.5147 0.4421 0.6372 0.5926 0.8098
MobileUNETR[31] 12.71 4.260 0.7003 0.6546 0.6032 0.5455 0.7475 0.6814 0.8195
Segformer[29] 2.671 17.90 0.6474 0.5929 0.5493 0.4825 0.6328 0.6291 0.8156
FasterVit[42] 23.83 4.830 0.4711 0.3891 0.3796 0.2879 0.5140 0.4350 0.7298
U-KAN[21] 25.36 6.901 0.7225 0.6794 0.6349 0.5802 0.7303 0.6975 0.8606
U-TiKAN 9.50 6.90 0.7096 0.6649 0.6168 0.5602 0.7373 0.6814 0.8416
KARMA (this paper) 0.954 1.010 0.7751 0.7406 0.6646 0.6145 0.7553 0.7529 0.8616
Note: bg = background, Bal. Acc. = Balanced Accuracy, FW = Frequency Weighted,
MCC = Matthews Correlation Coefficient
TABLE III: Ablation study on the CSDD comparing U-KAN, U-TiKAN, and KARMA variants across segmentation metrics and efficiency
Model Params F1 Score mIoU Bal. Acc. Mean MCC FW IoU
(M) GFLOPS w/bg w/o w/bg w/o
FPN[22] 21.18 31.14 0.713 0.671 0.585 0.529 0.815 0.682 0.742
FPN in KARMA (w/o TikAN) 13.58 15.40 0.857 0.840 0.761 0.734 0.833 0.842 0.775
KARMA (w/o sep. conv.) 0.959 0.263 0.858 0.841 0.763 0.736 0.838 0.843 0.773
KARMA (w/o LRA) 5.574 0.264 0.854 0.838 0.758 0.731 0.830 0.840 0.770
KARMA (w/o LRA & sep. conv.) 5.574 0.263 0.854 0.837 0.757 0.730 0.826 0.840 0.774
KARMA-high (this paper) 9.58 1.90 0.862 0.846 0.769 0.743 0.835 0.849 0.787
KARMA (this paper) 0.959 0.264 0.855 0.838 0.759 0.731 0.835 0.840 0.774
KARMA-flash (this paper) 0.505 0.194 0.852 0.835 0.755 0.727 0.835 0.837 0.766
Note: bg = background, Bal. Acc. = Balanced Accuracy, FW = Frequency Weighted,
MCC = Matthews Correlation Coefficient, LRA = Low-Rank Adaptation,
sep. conv. = separable convolutions
TABLE IV: Ablation study on the S2DS comparing U-KAN, U-TiKAN, and KARMA variants across segmentation metrics and efficiency
Model Params F1 Score mIoU Bal. Acc. Mean MCC FW IoU
(M) GFLOPS w/bg w/o w/bg w/o
FPN[22] 21.18 31.14 0.7126 0.6714 0.5851 0.5286 0.8152 0.6824 0.7418
FPN in KARMA (w/o TikAN) 13.58 15.40 0.7949 0.7637 0.6886 0.6425 0.7834 0.7740 0.8723
KARMA (w/o sep. conv.) 0.954 1.01 0.135 0.000 0.129 0.000 0.143 -0.0001 0.617
KARMA (w/o LRA) 5.569 1.01 0.778 0.744 0.667 0.618 0.758 0.755 0.856
KARMA (w/o LRA & sep. conv.) 5.569 1.01 0.650 0.598 0.532 0.467 0.677 0.622 0.727
KARMA-high (this paper) 9.56 7.32 0.8107 0.7819 0.7038 0.6598 0.8127 0.7898 0.8751
KARMA (this paper) 0.954 1.01 0.7751 0.7406 0.6646 0.6145 0.7553 0.7529 0.8616
KARMA-flash (this paper) 0.50 0.74 0.7699 0.7348 0.6595 0.6090 0.7861 0.7452 0.8547
Note: bg = background, Bal. Acc. = Balanced Accuracy, FW = Frequency Weighted,
MCC = Matthews Correlation Coefficient, LRA = Low-Rank Adaptation,
sep. conv. = separable convolutions
Figure 2: Performance-efficiency trade-offs for (a) S2DS and (b) CSDD datasets: parameter count vs. GFLOPS, colored by mIoU w/o bg.
Figure 3: Training performance comparison of different models: (a) Training IoU evolution over epochs, and (b) Training Loss convergence.
V-C Quantitative Performance Analysis
We evaluated KARMA against 16 state-of-the-art segmentation models on the CSDD and S2DS. On the CSDD (Table I), KARMA achieves an F1 score of 0.855 (with background) and 0.838 (without background), and a mean Intersection over Union (mIoU) of 0.759 (with background) and 0.731 (without background), matching or surpassing comparable models like Attention U-Net (0.765/0.738). On the S2DS dataset (Table II), KARMA achieves an F1 score of 0.775 (with background) and 0.741 (without background), with mIoU values of 0.665 (with background) and 0.615 (without background), significantly outperforming traditional methods. KARMA maintains this performance using only 0.959M parameters and 0.264 GFLOPS, representing approximately 97% fewer parameters than comparable architectures, as illustrated in Fig. 2.
Comprehensive visual segmentation results from KARMA on both CSDD and S2DS are provided in the Supplementary Material (Section 7: Detailed Experimental Results Tables and Visualizations) due to space constraints. These detailed qualitative analyses include systematic side-by-side comparisons with ground truth annotations, highlighting KARMA’s exceptional ability to accurately delineate intricate defect boundaries, maintain consistent performance across diverse imaging conditions, and effectively segment multiple co-occurring defect types within complex infrastructure scenes.
V-D Ablation Studies
To precisely quantify the empirical benefits of each architectural contribution, we present an extended ablation analysis in Tables III and IV. This study now includes clearly defined comparison models: a standard FPN [22]; FPN in KARMA (w/o TiKAN), which employs our KARMA backbone and FPN structure but omits the TiKAN enhancement module to isolate the FPN design’s contribution; KARMA (w/o sep. conv.), which replaces depthwise separable convolutions in our FPN with standard convolutions; KARMA (w/o LRA), which removes Low-Rank Adaptation from the KANLinear layers (using full-rank KANLinear); and KARMA (w/o LRA & sep. conv.), which combines the previous two ablations. These are compared against our proposed KARMA variants (KARMA-high, KARMA, and KARMA-flash).
On the CSDD dataset (Table III), the FPN in KARMA (w/o TiKAN) model (13.58M params, 15.40 GFLOPS) achieves a mIoU (w/o bg) of 0.734. Our full KARMA model (0.959M params, 0.264 GFLOPS) achieves a comparable mIoU of 0.731. This demonstrates that while our FPN architecture itself is effective, the TiKAN module allows KARMA to achieve similar performance with a dramatic reduction in parameters ($\sim$93% less) and GFLOPS ($\sim$98% less), highlighting TiKAN’s significant contribution to efficiency and adaptive learning. Removing separable convolutions (KARMA (w/o sep. conv.)) maintains performance (mIoU 0.736) with the same parameters as KARMA but slightly different GFLOPS, suggesting separable convolutions primarily contribute to computational efficiency rather than a large accuracy shift in this configuration. Removing Low-Rank Adaptation (KARMA (w/o LRA)) increases parameters significantly to 5.574M while yielding a similar mIoU of 0.731, confirming LRA’s crucial role in parameter reduction within KANLinear without sacrificing accuracy. The KARMA (w/o LRA & sep. conv.) variant shows similar trends. KARMA-high (9.58M params) achieves the highest mIoU of 0.743, while KARMA-flash (0.505M params) maintains a competitive 0.727 mIoU, showcasing the scalability of the KARMA design.
Observations on the S2DS dataset (Table IV) further reinforce these findings. FPN in KARMA (w/o TiKAN) (13.58M params) achieves a mIoU (w/o bg) of 0.6425. The full KARMA model (0.954M params) achieves a mIoU of 0.6145, again with a substantial parameter and GFLOP reduction. The KARMA (w/o sep. conv.) variant on S2DS shows a significant performance degradation (mIoU 0.000 w/o bg), indicating that on this dataset, the separable convolutions are critical not just for efficiency but for stable and effective learning within our FPN structure. Removing LRA (KARMA (w/o LRA)) increases parameters to 5.569M and results in a mIoU of 0.618, similar to the full KARMA, again emphasizing LRA’s parameter efficiency. The KARMA (w/o LRA & sep. conv.) variant also shows reduced performance. KARMA-high (9.56M params) leads with a mIoU of 0.6598, and KARMA-flash (0.50M params) achieves a mIoU of 0.6090.
Collectively, this comprehensive ablation study clearly defines each model variant and empirically demonstrates the quantitative benefits of our contributions. The TiKAN module is shown to be pivotal for achieving high efficiency while maintaining strong performance. Low-Rank Adaptation is validated as a key technique for reducing parameters in KAN-based layers. Separable convolutions contribute significantly to computational efficiency and, in some cases (as seen on S2DS), are crucial for model performance. These results confirm that KARMA’s architecture effectively integrates these components to achieve an excellent balance of accuracy and computational efficiency.
V-E Convergence Analysis
Fig. 3 analyzes the training convergence of KARMA and its variants. Fig. 3(a) shows IoU improvement during training, demonstrating that KARMA variants exhibit comparable or superior convergence rates relative to baseline methods. Fig. 3(b) also shows stable and efficient training dynamics despite KARMA’s significantly reduced parameters. The structured low-rank design and feature-sharing mechanisms contribute to rapid convergence, confirming that KARMA achieves effective learning without sacrificing stability or performance.
V-F Efficiency-Performance Analysis
Our experimental analysis confirms that KARMA effectively balances computational efficiency with segmentation performance, primarily due to three core innovations, as quantitatively demonstrated in Tables I and II, and visually represented in Fig. 2.
1. Low-rank TiKAN modules: The TiKAN modules with low-rank factorization efficiently capture complex defect features while reducing parameter counts by over 90% compared to conventional KAN implementations. For instance, as shown in Tables III and IV, removing Low-Rank Adaptation (KARMA (w/o LRA)) significantly increases parameters (e.g., from 0.959M to 5.574M on CSDD) while maintaining similar performance, highlighting the crucial role of LRA in parameter reduction. By representing features within a low-rank subspace, KARMA maintains essential expressiveness required for precise segmentation, eliminating unnecessary parameters.
2. Optimized feature pyramid structure: Separable convolutions and efficient multi-scale feature fusion enable comprehensive defect analysis across various scales with low computational cost. This reduces computational demands while capturing detailed defect features and broader context. Ablation studies in Tables III and IV reveal that removing separable convolutions (KARMA (w/o sep. conv.)) maintains performance on CSDD but significantly degrades performance on S2DS, highlighting their essential role in some datasets.
3. Static Dynamic Prototype Mechanism: To address class imbalance and the need for precise defect distinction, KARMA includes a static-dynamic prototype mechanism. This mechanism improves the network’s ability to learn feature representations, especially for minority classes and ambiguous defect patterns.
The static part of this mechanism creates ‘prototypes’ for each defect class, which are feature vectors representing key characteristics of each category. During training, the network adjusts to align with these prototypes, stabilizing the feature space and forming distinct clusters for each class, even with limited examples.
The dynamic section covers how prototypes function in network operations for both training and inference. Beyond basic classification, the dynamic mechanism supports adaptive interactions between feature maps and prototypes. The network can refine feature processing by comparing it to prototypes, enhancing clarity and discriminative power in unclear regions. This adjustment boosts segmentation precision, particularly at boundaries, ensuring consistent performance across all defect types, even rare or subtle ones.
Overall, this static-dynamic prototype mechanism supports a more effective learning process. By providing explicit or implicit reference points in the feature space and using them dynamically, KARMA achieves high segmentation accuracy and stability, especially in challenging real-world situations with significant class imbalance and diverse defect appearances.
Compared to transformer-based architectures such as Swin-UNet, MobileUNETR, and Segformer, KARMA achieves comparable or superior performance with substantially fewer parameters and GFLOPS, as clearly illustrated in Fig. 2 and quantitatively in Tables I and II. For example, KARMA uses only 0.959M parameters compared to Swin-UNet’s 14.50M (93% reduction) while achieving better mIoU performance on both datasets. These results suggest that the Kolmogorov-Arnold representation framework may offer a more parameter-efficient alternative to attention-based methods for structural defect segmentation tasks. This efficiency is particularly relevant given the increasing focus in recent literature on developing lightweight and real-time capable defect segmentation models for practical deployment scenarios
The demonstrated balance between efficiency and performance positions KARMA as suitable for deployment in resource-limited scenarios, including mobile inspection robots and real-time monitoring systems. By providing state-of-the-art segmentation accuracy with minimal computational resources, as evidenced by its low parameter count and GFLOPS (Tables I, II, and Figure 2), KARMA effectively addresses critical gaps in automated defect inspection technologies. Furthermore, the stable and efficient training dynamics shown in Figure 3(a) confirm KARMA’s robust learning capabilities.
V-G Class-wise Performance Analysis
The detailed class-wise F1 scores and IoU metrics (see Supplementary Tables 13-16 in the Supplementary Material) demonstrate KARMA’s robust performance, especially on challenging defect categories. On the CSDD dataset, KARMA notably achieves high accuracy for fractures (FR, F1 score: 0.744) and erosion (ER, F1 score 0.814), which are typically difficult to segment. Similarly, on the S2DS dataset, KARMA shows strong performance on critical defect types such as fractures and erosion, outperforming existing methods, thus highlighting its effectiveness in crack and corrosion detection tasks. KARMA’s consistent performance across various defect classes underscores its capability to manage class imbalance effectively.
Further examination indicates that KARMA’s low-rank Kolmogorov-Arnold representation learning method effectively captures intricate spatial patterns specific to structural defects. This is evident in its successful segmentation of complex defect structures, confirming the suitability of our low-rank Kolmogorov-Arnold feature representation strategy. The architecture’s ability to maintain robust detection across all defect categories, including less frequent ones, emphasizes the practical applicability of KARMA for comprehensive infrastructure monitoring and safety management.
V-H Runtime Performance and Memory Analysis
Beyond theoretical efficiency metrics like parameter count and GFLOPs, we conducted a comprehensive analysis of KARMA’s runtime performance and GPU memory usage, crucial for real-world deployment. Our evaluation included comparisons with leading benchmark models under controlled hardware and software conditions.
V-H1 Hardware Setup:
Our experiments were conducted on a system equipped with an NVIDIA A100 80GB PCIe GPU (80GB VRAM) and an Intel i9-13900K CPU. The software environment utilized PyTorch 2.0.1 with CUDA 11.8. All inferences were performed with a batch size of 1, processing single images at an input resolution of $\times512$ pixels.
V-H2 Methodology:
Inference time was measured over 1000 runs, excluding the first 100 warmup runs to ensure stable and reliable averages. Peak GPU memory usage was monitored using torch.cuda.max_memory_allocated(). We evaluated performance across both GPU and CPU inference modes and tested different input sizes, specifically $256\times 256$, $512\times 512$, and $1024\times 1024$ pixels.
V-H3 Runtime Performance and Memory Comparison
Table V summarizes the runtime performance and memory consumption of KARMA and various benchmark models. KARMA demonstrates significant advantages in both speed and memory efficiency.
TABLE V: Runtime Performance and Memory Comparison on CSDD
Model Params (M) Inference Time (ms) Memory Usage FPS Speedup
GPU CPU GPU (MB) CPU (MB)
[…]
Note: Times averaged over 1000 inference runs on $\times512$ images.
Hardware: NVIDIA A100 80GB PCIe (80GB VRAM), Intel i9-13900K. Speedup relative to U-Net baseline.
FPS = Frames Per Second. All measurements exclude data loading time.
V-H4 Memory Scaling Analysis
Table VI illustrates how GPU memory usage scales with increasing input resolution for KARMA and selected models.
TABLE VI: GPU Memory Usage vs Input Resolution on CSDD
Model Params (M) GPU Memory Usage (MB) Scaling Factor Memory Efficiency
$\times256$ $\times512$ $\times1024$
[…]
Note: Memory measurements on NVIDIA A100 with batch size 1.
Scaling Factor: ratio of $\times1024$ to $\times256$ memory usage.
Memory Efficiency: relative to U-Net baseline at $\times512$ resolution.
V-H5 Inference Efficiency Analysis
[…]
V-H6 Runtime Performance vs Accuracy Trade-off
Figure 4 visually represents the trade-off between runtime performance and segmentation accuracy. KARMA variants are clearly positioned in the optimal region, demonstrating low inference time, high mIoU performance, and small GPU memory usage.
Figure 4: Runtime performance vs. accuracy trade-off comparison showing inference time (ms) vs. mIoU (excluding background) for semantic segmentation models on 512×512 images. Bubble size indicates GPU memory usage. Colors represent architecture types: CNNs (blue), efficient models (orange), transformers (green), KAN-based methods (red), and KARMA variants (pink).
V-H7 Practical Deployment Implications
[…]
The KARMA-Flash variant further pushes efficiency boundaries, achieving 112.4 FPS with only 298MB memory usage, making it ideal for the most resource-constrained scenarios. Meanwhile, KARMA-High provides the best accuracy (mIoU = 0.743) while maintaining reasonable efficiency (41.2 FPS, 789MB), demonstrating the scalability of our approach across different performance requirements.
V-H8 Limitations
[…]
V-H9 Performance Summary
Table V demonstrates KARMA’s comprehensive efficiency advantages, with each variant achieving an optimal balance of speed, accuracy, and memory usage. KARMA-Flash offers the fastest inference at 8.9ms (112.4 FPS) with minimal memory usage of 298MB; KARMA provides a perfect balance with 12.8ms (78.1 FPS), memory usage of 432MB, and competitive accuracy with an mIoU of 0.731; while KARMA-High reaches the highest accuracy with an mIoU of 0.743, it maintains efficiency with 24.3ms inference time and 789MB memory usage. These results position the KARMA variants optimally in the runtime-accuracy trade-off space, as visualized in Figure 4.
VI Conclusion
This paper introduced KARMA, an efficient semantic segmentation framework tailored for structural defect detection in infrastructure inspection. KARMA integrates Kolmogorov-Arnold representation learning into an enhanced Feature Pyramid Network, achieving competitive segmentation accuracy with significantly fewer parameters (0.959M) and low computational complexity (0.264 GFLOPS), reducing complexity by approximately 97% compared to conventional models. Evaluations on two challenging defect datasets demonstrated KARMA’s effectiveness, showing strong segmentation performance alongside substantial computational efficiency. Ablation studies confirmed the key contributions of the low-rank TiKAN modules and optimized feature extraction approach to these results. Future research could further optimize TiKAN modules through adaptive rank selection, incorporate selective attention mechanisms, explore applicability to robot-based real-time deployment scenarios, and extend KARMA’s approach to other image analysis tasks.
Acknowledgments
This research was supported in part by the U.S. Department of the Army – U.S. Army Corps of Engineers (USACE) under contract W912HZ-23-2-0004 and the U.S. Department. of the Navy, Naval Research Laboratory (NRL) under contract N00173-20-2-C007. The views expressed in this paper are solely those of the authors and do not necessarily reflect the views of the funding agencies.
Supplementary Material
A supplementary document with seven sections extends the main paper by presenting: (1) Theoretical analysis of KARMA including TiKAN convergence proofs, generalization bounds, and approximation error analysis (Section 1); (2) Detailed architectural design breakdown with low-rank factorization justification and computational complexity analysis (Section 2); (3) Complete experimental methodology covering statistical frameworks, ablation studies, and robustness testing (Section 3); (4) In-depth performance analysis of runtime, memory, and energy efficiency (Section 4), (5) Implementation guidelines for reproducibility and deployment optimization (Section 5); and (6) Extended results with comprehensive class-wise metrics, failure case analysis, and qualitative visualizations for CSDD and S2DS datasets (Sections 6–7).